Determine the margin of error for a confidence interval to estimate the population mean with n=18 and s​ =14.5 for the confidence levels below.

​a)

808%

​b)

90%

​c)

99%

​a) The margin of error for an 80​% confidence interval is

​(Round to two decimal places as​ needed.)
Answer #1

(a)

n = Sample Size = 18

s = 14.5

SE= s/\sqrt{n}

= 14.5/\sqrt{18} = 3.4177

ndf = 18- 1 =17

\alpha = 0.20

From Table critical values of t = \pm 1.3334

So,

Margin of Error = t X SE

= 1.3337 X 3.4177 = 4.56

(b)

n = Sample Size = 18

s = 14.5

SE= s/\sqrt{n}

= 14.5/\sqrt{18} = 3.4177

ndf = 18- 1 =17

\alpha = 0.10

From Table critical values of t = \pm 1.7396

So,

Margin of Error = t X SE

= 1.7396 X 3.4177 = 5.95

(c)

n = Sample Size = 18

s = 14.5

SE= s/\sqrt{n}

= 14.5/\sqrt{18} = 3.4177

ndf = 18- 1 =17

\alpha = 0.01

From Table critical values of t = \pm 2.8982

So,

Margin of Error = t X SE

= 2.8982 X 3.4177 = 9.91

Thus, we get:

(a) 80 % 4.56
(b) 90 % 5.95
(c) 99% 9.91

(a) The margin of error for an 80% confidence interval is 4.56